Calculate the following salespeople’s gross wages

Employee        Sales            Returns            Commission      Draw               Quota

                                                                        Rate                

A. Marge         $40,000            0                      15%                  0                       0

B. George        $38,000           0                      13%                 $600                0

C. Joan            $81,000           0                      10%                 $1,000             $20,000

D. Steve          $52,300       $1,400                  9.5%                0                    $10,000
Answer #1

a) Marge Gross Wages = 40000*15% = 6000
b) George Gross Wages = 38000* 13% - Draw = 4940-600 = 4340
c) Joan Gross wages = 81000*10%-Draw + Quota = 8100 -1000 + 20000 = 28000
d) Steve Gross Wages = (52300-1400)*9.5% + 10000 = 14835.50

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well